I am having trouble with the story problems

Three consectutive even integers are such that the square of the third is 76 more than the square of the second, Find the three integers.

Answer by checkley77(12844)

Let x, x+2 & x+4 be the 3 even integers.
(x+4)^2=(x+2)^2+76
x^2+8x+16=x^2+4x+4+76
x^2-x^2+8x-4x=-16+4+76
4x+64
x=64/4
x=16 ans. for the smallest integer.
16+2=18 ans. for the middle integer.
16+4=20 ans. for the largest integer.
Proof:
20^2=18^2+76
400=324+76
400=400